OpenStudy (brittany16):

Solve the equation. -11=18+s/5

OpenStudy (oaktree):

-145

OpenStudy (anonymous):

-29= s/5 -145=s

OpenStudy (anonymous):

-11 -18 = s/5 -29 = s/5 -29 * 5 = s -145 = s

OpenStudy (anonymous):

Christ on a biscuit! Put these into Wolfram Alpha! http://i.imgur.com/vTBXQ.png